Math 142  University of South Carolina
D. Meade  Spring 1996

Math 142 (Section 502) – Calculus II

Instructor Professor Doug Meade

Office Hours: MW 10:00–11:00, and by prior appointment
Office:  LeConte College 300E
Phone:  777-6183

E-mail:  meade@math.sc.edu

WWW URL  http://www.math.sc.edu/~meade/math142/

Meeting Times MW 11:15am–12:05pm, LC 303B TTh 11:00am–12:15pm, LC 303B

Text Varberg and Purcell, Calculus with Analytic Geometry.

Prerequisite Completion of Math 141 with a grade of C or better or qualification through placement.

Course Content This is a continuation of Math 141. A working understanding of all the major topics from Math 141: limits, differentiation, integration, extrema, Fundamental Theorem of Calculus, etc. is essential for success in this course. Almost all topics from Chapters 7 – 12 will be discussed in this course. Some of the specific topics that will be covered include: Chapter 7 Transcendental Functions

• logarithms

• inverse functions Chapter 8 Techniques of Integration

• substitution

• integration by parts

• partial fractions Chapter 9 Indeterminate Forms and Improper Integrals

• l’Hôpital’s Rule

• integration when the Fundamental Theorem does not apply Chapter 10 Numerical Methods and Approximations

• approximation of functions

• numerical integration Chapter 11 Sequences and Series

• convergence tests

• power series Chapter 12 Polar Coordinates

• graphing

• area Study Hints Reading the material in advance of the lecture is strongly encouraged. Benefits of this

preparation include obtaining a familiarity with the terminology and concepts that will be encountered (so you can distinguish major points from side issues), being able to formulate questions about the parts of the presentation that you do not understand, and having a chance to review the skills and techniques that will be needed to apply the new concepts.

For additional assistance, do not forget about the Math Lab located in LeConte College 101. Assistance for all 100- and 200-level courses is available on from 1:00–8:00 MTWTh and from 1:00–3:00 on F.

Finally, as previously mentioned, you are assumed to have a mastery of the topics in Math 141. If you are not comfortable with your basic Calculus skills, please discuss your concerns with me before they become a problem.

Grading Your grade in this course will be based on your performance on quizzes, four (4) mid-term exams, and a final exam. The weights assigned to each of these components will be:

Quizzes  10%
Mid-term exams (3) 60%
Final exam  30%

Course grades will be determined according to the following scale:

A 90 –100
B 80 – 89
C 70 – 79
D 60 – 69
F 0 – 59

Note that the deadline to drop this course with a grade of W is Friday, \underline{February 16, 1996}.

Exams The lowest of your four (4) mid-semester exam scores will not be used in determining your overall grade. Tentative dates for these exams are:

Thursday, January 25  Chapter 7
Thursday, February 15  Chapter 8 (and part of 9)
Thursday, March 21  Chapters 9, 10 (and part of 11)
Thursday, April 18  Chapter 11

Make-up exams will be given only for documented reasons of illness, family emergency or participation in a University sponsored event. Excuses such as oversleeping, forgetting the time or location of the exam, and lack of studying are explicitly noted as unacceptable grounds for the administration of a make-up exam.

A comprehensive final will be given at \underline{9:00a.m.} on \underline{Monday, April 29, 1996}. (Do not trust exam schedules published by unofficial sources, e.g. the Mortar Board.)

Homework Problems will be assigned on a regular basis. You are expected to work all of these problems. We will go over some, but not all, of these problems in class. Your solutions will not be collected.

Quizzes There will be a quiz each Monday of the semester. Each quiz will consist of one or two problems similar in nature to the homework problems. Your quiz grade will be computed by your ten (10) highest quiz scores. No make-up quizzes will be permitted.

Academic Honesty Cheating and plagiarism will not be tolerated in this course. You may discuss homework problems with others, but do not copy solutions from another student or from a book. Violations of this policy will be dealt with in a matter consistent with University guidelines.